OpenStudy (anonymous):

Can someone tell me if I did this right. Simplify. Assume that all nonnegative real numbers Square root of 50

OpenStudy (anonymous):

\[\sqrt{50}\]

OpenStudy (anonymous):

\[=\sqrt{10}\times \sqrt{5}\]

OpenStudy (anonymous):